# BugCraft: End-to-End Crash Bug Reproduction Using LLM Agents in Minecraft

### v1

@@
Abstract.
Context: Reproducing game bugs, in our case crash bugs in continuously evolving games like Minecraft, is a notoriously manual, time-consuming, and challenging process to automate. Despite the success of LLM-driven bug reproduction in other software domains, games, with their complex interactive environments, remain largely unaddressed.
Objective: This paper introduces BugCraft, a novel end-to-end framework designed to automate the reproduction of crash bugs in Minecraft directly from user-submitted bug reports, addressing the critical gap in automated game bug reproduction.
Method: BugCraft employs a two-stage approach: first, a Step Synthesizer leverages LLMs and Minecraft Wiki knowledge to transform bug reports into high-quality, structured steps to reproduce (S2R). Second, an Action Model, powered by a vision-based LLM agent (GPT-4o) and a custom macro API, executes these S2R steps within Minecraft to trigger the reported crash. To facilitate evaluation, we introduce BugCraft-Bench, a curated dataset of Minecraft crash bug reports.
Results: Evaluated on BugCraft-Bench, our framework successfully reproduced 30.23% of crash bugs end-to-end. The Step Synthesizer demonstrated a 66.28% accuracy in generating correct bug reproduction plans, highlighting its effectiveness in interpreting and structuring bug report information.
Conclusion: BugCraft demonstrates the feasibility of automated reproduction of crash bugs in complex game environments using LLMs, opening promising avenues for game testing and development. The framework and the BugCraft-Bench dataset pave the way for future research in automated game bug analysis and hold potential for generalization to other interactive game platforms. Finally, we make our code open at https://bugcraft2025.github.io/
Automated Bug Reproduction, Vision-Based Agent, Large Language Model, Crash Bugs, Minecraft, Bug Tracking
copyright: ccjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/24/06ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Maintaining softwareccs: Software and its engineering Interactive games
1. Introduction
@@
In the maintenance phase of the game development cycle, it is crucial to provide updates that improve the game to maintain player retention and enhance engagement (Zhong_Xu_2021, ). These game updates should also fix bugs for improved user experience. As a type of bug, crash bugs are especially severe because they prevent the user from experiencing the game. They are also difficult to fix, being the most recurring bug type even after updates addressing the bug (Truelove_Santana_De_Almeida_Ahmed_2021, ).
@@
A primary source of information for these updates is the user-submitted bug reports. A prime example is Mojira [^1] https://bugs.mojang.com/issues, the public bug tracker for Minecraft developer Mojang Studios. Mojira hosts more than 450,000 bug reports, with over half related to Minecraft: Java Edition [^2] https://www.minecraft.net/en-us/about-minecraft. However, many of these reports are of low quality, often missing essential components like steps to reproduce (S2R), expected outcomes, and can have mistakes in them (Zimmermann_Premraj_Bettenburg_Just_Schroter_Weiss_2010, ; Soltani_Hermans_Bäck_2020, ). The studies showing low-quality bug reports across numerous projects suggest that Mojira could potentially face similar challenges.
@@
Due to the excessive amount of labor involved, game testing is an entire field in itself, and it is a notoriously manual process that takes a significant amount of time (Politowski_Petrillo_Gueheneuc_2021, ). Most games lack automated test frameworks because implementing and generalizing such systems is challenging. For example, Minecraft’s internal unit test framework GameTest [^3] https://www.minecraft.net/en-us/creator/article/get-started-gametest-framework lacks the tools to reproduce UI movement and precise player interactions, only covering a part of what is a huge sandbox game.
@@
Looking at a more general usage, automated bug reproduction systems have been developed (Kang_Yoon_Yoo_2023, ). With the advent of large language models (LLMs) that are capable of tool calling (i.e., interacting with external tools/APIs) (qin2024toollearningfoundationmodels, ) and code generation (jiang2024surveylargelanguagemodels, ), the implementations of these systems have become more general. Similar works (Wang_Zhao_Feng_Zhang_Halfond, ) exist for reproducing bugs in Android apps through interacting with the available GUI. These works are limited in interacting with the game, which needs action taking in a complex environment. To our knowledge, there are no works for automating bug reproduction in an open-ended game environment like Minecraft.
In this paper, we present BugCraft, an agentic bug reproduction framework for Minecraft. BugCraft can take in the initial bug report, synthesize high-quality steps to reproduce bugs alongside relevant game data from external sources and reproduce them step by step through scripting. In the end, our framework generates a log that can reproduce the bug. In addition, due to the lack of a dataset of crash bugs in a game, we present BugCraft-Bench, a crash bug dataset collected through Mojira. Our dataset contains 86 manually confirmed, reproducible crash bug reports. Finally, our reproduction package, alongside with our results can be found at: https://figshare.com/s/70b5d141e69ff822cd40.
@@
Figure 1. The BugCraft framework, illustrating the two-stage process of S2R synthesis and action model execution.
Our paper begins by discussing related work in Section 2, followed by an overview of the framework in Section 3. We define our research questions in Section 3.3, evaluate the framework in Section 4, and discuss our findings in Section 5. The paper mentions the threats to validity in Section 6, and concludes with a summary of contributions in Section 7.
2. Related Work
This section provides an overview of related topics, focusing on bug reproduction in games and the development of agents powered by LLMs. We also discuss similar research efforts relevant to our framework.
2.1. Bug Reproduction in Games
@@
Game testing is traditionally a manual and time-consuming process, and the primary approach involves playtesting, in which testers manually play the game to identify issues (Politowski_Petrillo_Gueheneuc_2021, ). Automated unit testing is often challenging for game developers due to the dynamic nature of game code and the complexity involved in capturing game states programmatically (Pascarella_Palomba_Di_Penta_Bacchelli_2018, ). In continuous development cycles, it becomes essential for developers to address existing bugs through regular updates to improve the gameplay experience. However, certain types of bugs, particularly crash bugs, are prone to reoccurrence even after an update is released to fix them (Truelove_Santana_De_Almeida_Ahmed_2021, ). One developer in a recent survey attributed this issue to limited or non-available testing time, highlighting the constraints of manual testing in the fast-paced, deadline-driven environment typical of game development.
Wuji (Zheng_Xie_Su_Ma_Hao_Meng_Liu_Shen_Chen_Fan_2019, ) represents the closest related work in automated game testing, using deep reinforcement learning to explore the game model to discover new bugs. However, Wuji’s capabilities are limited to predefined game areas and narrow action spaces. It requires manual environment setup without the ability to follow specific test prompts. Furthermore, Wuji operates in the domain of bug finding, while our framework focuses on reproducing known bugs. Our approach provides end-to-end bug reproduction capabilities, starting from game initialization and covering specific scenarios, enabling a more comprehensive approach to addressing reported issues.
2.2. Automated Bug Reproduction Using LLMs
@@
Automated bug reproduction entails converting unstructured bug reports into executable scripts reliably replicating the reported issues. LLMs have shown considerable success in generating code, particularly in creating unit tests (Kang_Yoon_Yoo_2023, ) and enhancing bug report details (Bo_Ji_Sun_Zhang_Wu_Wei_2024, ). Recent frameworks leveraging LLMs have been developed to automate bug reproduction in mobile applications, particularly within Android environments, by interacting with the GUI (Wang_Zhao_Feng_Zhang_Halfond, ). These frameworks enable a series of structured actions where the LLM makes decisions based on the bug report, GUI interactions, and prior actions while additional contextual information is provided as needed by the framework. Frameworks in (Wang_Zhao_Feng_Zhang_Halfond, ) and (Feng_Chen_2024, ) rely on methods that convert the GUI into textual representations, allowing the LLM to interact with and manipulate the interface through text-based commands. Additionally, a study (Liu_Li_Chen_Wang_Wu_Wang_Hu_Wang_2024, ) incorporates visual inputs, enhancing the LLM’s capacity to interpret and interact with complex GUIs. This approach has achieved performance comparable to other vision-based methods, highlighting the potential for LLMs to effectively combine textual and visual data in automated bug reproduction tasks.
2.3. LLM Agents
@@
LLMs have shown great success at acting as artificial agents, which are entities capable of perceiving their surroundings, making decisions, and taking actions in response (xi2023rise, ; yao2023react, ). The agent framework involves equipping LLMs with tools, often in the form of APIs, to expand their action space and enable them to interact effectively with the environment. The most important part is that the LLMs were able to do this mainly by utilizing in-context learning (tombrownfewshot, ), alleviating the need to train specialized models for each in-domain task. Moreover, LLMs have also shown to be successful at taking action in an open-ended environment. This ability to act extends even to games like Minecraft, where the Voyager framework leveraged an LLM agent to achieve in-game progress milestones significantly faster than previous methods (wang2023voyager, ).
3. Methodology
Our experiment utilizes a custom dataset of Minecraft crash bug reports, processed through a custom framework called BugCraft, which operates in two main phases: creating the Steps to Reproduce (S2R) and executing the S2R (step clusters). Figure 3 provides a high-level overview of our framework. The following sections will elaborate on its components: we begin with the dataset creation process (subsection 3.1), followed by a brief introduction to Bugcraft (subsection 3.2). We then describe the preprocessing of a bug report (subsection 3.2.1), the Step Synthesizer component (subsection 3.2.2), the Action Model component (subsection 3.2.3), the general implementation details of the entire framework (subsection 3.2.4) and finally the evaluation setup (subsection 3.3).
3.1. Data Curation
@@
3.1.1. BugCraft-Bench Dataset Curation
@@
We built a custom dataset consisting of 86 Minecraft: Java Edition crash bug reports to evaluate our framework. These reports were collected from Mojira, Minecraft’s official bug tracker, utilizing Jira’s REST API. We employed a finegrained query to retrieve all reports classified under the ”Crash” category. Moreover, for each report, we gathered all associated comments, attachments, and external links (e.g., YouTube links), which may contain relevant files or videos. These were identified using regular expressions (Regex).
Figure 2. Bug Report Codenamed MC-276621
@@
In Figure 2, we present a sample bug report, identified as MC-276621444https://bugs.mojang.com/browse/MC-276621, which was verified and resolved by Mojang in a subsequent snapshot. This report illustrates a bug that can be reproduced through a specific combination of commands and UI interactions. Additionally, another user, contributed steps-to-reproduce (S2R), which were initially missing from the report. Importantly, the BR in its final state lacks commands that would make reproducing the bug easier. Furthermore, though not relevant to this particular report, some bug reports, such as MC-275657555https://bugs.mojang.com/browse/MC-275657, contain an implicit assumption that a version-specific setting is enabled without explicitly stating it, creating ambiguity during reproduction.
@@
Each report in the dataset was categorized according to its ”Confirmation Status” label in Mojira, which includes four possible states: ”Confirmed,” ”Community Consensus,” ”Plausible,” and ”Unconfirmed.” We primarily focused on ”Confirmed” and ”Community Consensus” reports, as these have been verified by the game’s developer, Mojang Studios, making them reliable ground truth instances. To focus solely on singleplayer bugs, we excluded reports related to servers, multiplayer interactions, and those requiring external files loaded into the game (such as datapacks, worlds, or world setting JSONs).
@@
To ensure the quality and relevance of our dataset, we employed a multi-stage filtration process, focusing specifically on single-player crash bugs. We began by targeting the most recent 300 (limited the size of the dataset because of resource constraints) crash bug reports from Mojira’s database following our query results, spanning from the most recent, MC-277967666https://bugs.mojang.com/browse/MC-277967 (resolved on 05/Nov/24, affecting version 24w44a), to MC-145102777https://bugs.mojang.com/browse/MC-145102 (created on 19/Feb/19, affecting version 19w08b), covering a total of 70 different game versions. We then downloaded all associated data for each of these 300 reports, including comments, attachments, and external links. Next, we undertook a manual review, a crucial step that involved categorizing each report based on several criteria. Our focus was to isolate single-player crash bugs, so we excluded reports involving multiplayer aspects like servers or Realms (a subscription based server service) [^8] https://www.minecraft.net/en-us/realms, those requiring interactions between different game versions, and any that depended on external resources such as datapacks, custom worlds, resource packs, or specific world settings. We also removed reports related to issues outside the game itself or those that were machine-dependent. Furthermore, reports were discarded if they were considered low quality due to inadequate information or if the final version differed so much from the initial draft that it could be considered a new submission. This meticulous review resulted in a refined set of 141 reports. Due to resource constraints and the significant cost associated with the LLM API calls, we chose to evaluate only a portion of the full dataset. As such, from the initial pool of 141 reports, we randomly selected a subset of 100 to create our final dataset. This dataset was used for detailed analysis and the evaluation of our framework.
3.1.2. Minecraft Wiki Dataset
@@
We use the official Minecraft Wiki [^9] https://minecraft.wiki site as our knowledge base. Since the base of the site used MediaWiki, we are able to get the wiki’s text using the provided wiki API. We pulled every page in the wiki that was not a redirect page. We stored each page as a separate file and used fuzzy matching with the page titles for retrieval. Using the wiki as our knowledge base allows us to circumvent reliance on our underlying LLM’s pre-training data.
3.2. BugCraft Framework
Figure 3. BugCraft Framework

### v2

# Agents in the Sandbox: End-to-End Crash Bug Reproduction for Minecraft

@@
Abstract
Reproducing game bugs, particularly crash bugs in continuously evolving games like Minecraft, is a notoriously manual, time-consuming, and challenging process to automate; insights from a key decision maker from Minecraft we interviewed confirm this, highlighting that a substantial portion of crash reports necessitate manual scenario reconstruction. Despite the success of LLM-driven bug reproduction in other software domains, games, with their complex interactive environments, remain largely unaddressed. This paper introduces BugCraft, a novel end-to-end framework designed to automate the reproduction of crash bugs in Minecraft directly from user-submitted bug reports, addressing the critical gap in automated game bug reproduction. BugCraft employs a two-stage approach: first, a Step Synthesizer leverages LLMs and Minecraft Wiki knowledge to transform bug reports into high-quality, structured steps to reproduce (S2R). Second, an Action Model, powered by a vision-based LLM agent and a custom macro API, executes these S2R steps within Minecraft to trigger the reported crash. To facilitate evaluation, we introduce BugCraft-Bench, a curated dataset of Minecraft crash bug reports. On BugCraft-Bench, our framework end-to-end reproduced 34.9% of crash bugs with GPT-4.1, outperforming baseline computer-use models by 37%. BugCraft demonstrates the feasibility of automated reproduction of crash bugs in complex game environments using LLMs, opening promising avenues for game testing and development. Finally, we make our code open at https://bugcraft2025.github.io.
I Introduction
@@
In the maintenance phase of the game development cycle, it is crucial to provide updates that improve the game to maintain player retention and enhance engagement [1]. These game updates should also fix bugs for improved user experience. As a type of bug, crash bugs are especially severe because they prevent the user from experiencing the game. They are also difficult to fix, being the most recurring bug type even after updates addressing the bug [2].
@@
A primary source of information for these updates is the user-submitted bug reports. A prime example is Mojira [^1] https://bugs.mojang.com/issues, the public bug tracker for Minecraft developer Mojang Studios. Mojira hosts more than 12 years of bugs for various versions of Minecraft [^2] https://bugs.mojang.com/browse/MC-1. However, many of these reports are of low quality, often missing essential components like S2R, expected outcomes, and can have mistakes in them [3, 4, 5, 6, 7]. These difficulties prolong bug lifecycles [8], as challenging reproduction decreases reviewer motivation and delays confirmation. The studies showing low-quality bug reports across numerous projects suggest that Mojira could potentially face similar challenges.
@@
Due to the excessive amount of labor involved, game testing is an entire field in itself, and it is a notoriously manual process that takes a significant amount of time [9]. To confirm the real-world impact of this challenge, we conducted a semi-structured interview with a key decision maker from Minecraft in May 2025. This expert emphasized the critical importance of successful reproduction, noting that once a bug is reproducible, a significant portion of the diagnostic challenge is considered overcome, bringing considerable relief to the development team. The interview confirmed that their teams manually reconstruct scenarios for a significant volume of crash reports, highlighting the substantial manual effort involved and the potential utility of automated solutions like BugCraft. This critical need for automation, especially in a huge sandbox game, motivates our work. Most games lack automated test frameworks because implementing and generalizing such systems is challenging. For example, Minecraft’s internal unit test framework GameTest [^3] https://www.minecraft.net/en-us/creator/article/get-started-gametest-framework lacks the support to reproduce UI movement and precise player interactions, only covering a part of what is a huge sandbox game. Moreover, developers need to write unit tests manually for each specific bug to be integrated into the game.
@@
Looking at a more general usage, automated bug reproduction systems have been developed [10]. With the advent of large language models (LLMs) that are capable of tool calling (i.e., interacting with external tools/APIs) [11] and code generation [12], the implementations of these systems have become more general. Similar works [13, 14, 15] exist for reproducing crash bugs in Android apps through interacting with the available GUI. These works are limited in interacting with the game, which needs action-taking in a complex environment. To our knowledge, there are no works for end-to-end automated crash bug reproduction in an open-ended game environment like Minecraft.
In this paper, we present BugCraft, an agentic bug reproduction framework for Minecraft, which makes the following novel contributions:
  • The first end-to-end framework for automated crash bug reproduction in a complex, 3D game environment. BugCraft tackles the full pipeline from an unstructured bug report to the generation of a script and a final log that can reliably reproduce the crash.
  • A novel, multi-stage Step Synthesizer that transforms ambiguous bug reports into actionable plans by synthesizing information from diverse sources, including user comments and domain-specific knowledge bases. This is achieved through a process of RAG, intelligent analysis of the game state to find reliable solutions, and a final structuring phase using what we term Step Clusters, granting the agent crucial flexibility during execution.
  • BugCraft-Bench, the first publicly available benchmark for this task, containing 86 manually confirmed and reproducible crash bug reports to facilitate future research.
We evaluate BugCraft against competitive baselines including OpenAI’s Computer Use Agent (OpenAI-CUA) and UI-TARS-1.5-7B [16], demonstrating a 34.9% success rate that outperforms existing approaches. Additionally, we test BugCraft with multiple LLMs (GPT-4o and GPT-4.1) to demonstrate how model improvements directly enhance framework performance. To ensure full reproducibility, we provide the BugCraft-Bench, the source code, run logs, and annotations at our project website: https://bugcraft2025.github.io.
@@
Figure 1: The BugCraft framework, illustrating the two-stage process of S2R synthesis and action model execution.
II Related Work
II-A Bug Reproduction in Games
@@
Game testing is traditionally a manual and time-consuming process, and the primary approach involves playtesting, in which testers manually play the game to identify issues [9]. Automated unit testing is often challenging for game developers due to the dynamic nature of game code and the complexity involved in capturing game states programmatically [17]. In continuous development cycles, it becomes essential for developers to address existing bugs through regular updates to improve the gameplay experience. However, certain types of bugs, particularly crash bugs, are prone to reoccurrence even after an update is released to fix them. One developer in a recent survey attributed this issue to limited or non-available testing time, highlighting the constraints of manual testing in the fast-paced, deadline-driven environment typical of game development [2].
Wuji [18], a system for automated game testing, utilizes deep reinforcement learning with the goal of discovering new bugs. This contrasts with BugCraft’s objective of reproducing known bugs from user reports. Therefore, due to its focus on bug finding rather than reproduction, and its limitations like narrow action spaces and manual setup, Wuji is not a direct baseline for our work.
II-B Automated Bug Reproduction Using LLMs
@@
Automated bug reproduction entails converting unstructured bug reports into executable scripts reliably replicating the reported issues. LLMs have shown considerable success in generating code, particularly in creating unit tests [10] and enhancing bug report details [19]. Recent frameworks leveraging LLMs have been developed to automate bug reproduction in mobile applications, particularly within Android environments, by interacting with the GUI [14]. These frameworks enable a series of structured actions where the LLM makes decisions based on the bug report, GUI interactions, and prior actions, while additional contextual information is provided as needed by the framework. Frameworks in [14] and [15] rely on methods that convert the GUI into textual representations, allowing the LLM to interact with and manipulate the interface through text-based commands. Additionally, a study [13] incorporates visual inputs, enhancing the LLM’s capacity to interpret and interact with complex GUIs. This approach has achieved performance comparable to other vision-based methods, highlighting the potential for LLMs to effectively combine textual and visual data in automated bug reproduction tasks.
II-C LLM Agents
@@
LLMs have shown great success at acting as artificial agents, which are entities capable of perceiving their surroundings, making decisions, and taking actions in response [20, 21]. The agent framework involves equipping LLMs with tools, often in the form of APIs, to expand their action space and enable them to interact effectively with the environment. The most important part is that the LLMs were able to do this mainly by utilizing in-context learning [22], alleviating the need to train specialized models for each in-domain task. Moreover, LLMs have also shown to be successful at taking action in an open-ended environment. This ability to act extends even to games like Minecraft, where the Voyager framework leveraged an LLM agent to achieve in-game progress milestones significantly faster than previous methods [23].
III Methodology
We introduce BugCraft, a novel framework for end-to-end crash bug reproduction in complex interactive environments, using a custom dataset of Minecraft bug reports. As illustrated in Figure 3, BugCraft operates in two phases: creating and executing S2R. The following sections detail our dataset (subsection III-A), BugCraft’s components (subsections III-B1III-B3), implementation (subsection III-B4), and evaluation setup (Subsection III-C).
III-A Data Curation
@@
III-A1 BugCraft-Bench Dataset Curation
@@
We built a custom dataset consisting of 86 Minecraft: Java Edition crash bug reports to evaluate our framework. These reports were collected from Mojira, Minecraft’s official bug tracker, utilizing Jira’s REST API. We used a fine-grained query to retrieve all reports classified under the “Crash” category. Moreover, for each report, we gathered all associated comments, attachments, and external links (e.g., YouTube links), which may contain relevant files or videos. These were identified using regular expressions.
Figure 2: Bug Report Codenamed MC-276621
@@
In Figure 2, we present a sample bug report, identified as MC-276621444https://bugs.mojang.com/browse/MC-276621, which was verified and resolved by Mojang in a subsequent snapshot. This report illustrates a bug that can be reproduced through a specific combination of commands and UI interactions. Additionally, another user, contributed steps-to-reproduce (S2R), which were initially missing from the report. Importantly, the bug report (BR) in its final state lacks commands that would make reproducing the bug easier. Furthermore, though not relevant to this particular report, some bug reports, such as MC-275657555https://bugs.mojang.com/browse/MC-275657, contain an implicit assumption that a version-specific setting is enabled without explicitly stating it, creating ambiguity during reproduction.
@@
Each report in the dataset was categorized according to its ”Confirmation Status” label in Mojira, which includes four possible states: ”Confirmed,” ”Community Consensus,” ”Plausible,” and ”Unconfirmed.” We primarily focused on ”Confirmed” and ”Community Consensus” reports, as these have been verified by the game’s developer, Mojang Studios, making them reliable ground truth instances.
@@
To ensure the quality and relevance of our dataset, we employed a multi-stage filtration process, focusing specifically on single-player crash bugs. We began by targeting the most recent bugs from the last five years (limited the size of the dataset because of resource constraints) crash bug reports from Mojira’s database following our query results, spanning from the most recent, MC-277967666https://bugs.mojang.com/browse/MC-277967 (resolved on 05/Nov/24, affecting version 24w44a), to MC-145102777https://bugs.mojang.com/browse/MC-145102 (created on 19/Feb/19, affecting version 19w08b), covering a total of 70 different game versions. We then downloaded all associated data for each of these reports, including comments, attachments, and external links. Next, we undertook a manual review to categorize each report based on several criteria. Our focus was to isolate the single-player crash bugs, so we excluded reports involving multiplayer aspects, those requiring interactions between different game versions, and any that depended on external resources. We also removed reports related to issues outside the game itself and discarded reports if the final version differed so much from the initial draft that it could be considered a new submission. This review resulted in a refined set of 141 reports. Due to resource constraints and the significant cost associated with the LLM API calls, we chose to evaluate only a portion of the full dataset. As such, from the initial pool of 141 reports, we randomly selected a subset of 100 to create our final dataset. After our testing of the bug reports manually, we eliminated 14 of the bug reports due to being irreproducible, leaving us with a final number of 86 bug reports. This dataset was used for detailed analysis and the evaluation of our framework.
III-A2 Minecraft Wiki Dataset
@@
We use the official Minecraft Wiki [^8] https://minecraft.wiki site as our knowledge base. Since the base of the site used MediaWiki, we are able to get the wiki’s text using the provided wiki API. We pulled every page in the wiki that was not a redirect page. We stored each page as a separate file for retrieval augmented generation (RAG). Using the wiki as our knowledge base allows us to circumvent reliance on our underlying LLM’s pre-training data.
III-B BugCraft Framework
Figure 3: BugCraft Framework